OpenStudy (anonymous):

what is the derivative of (x^2)/(e^(4x))

OpenStudy (shinalcantara):

\[\frac{ x ^{2} }{ e ^{4x} }\] since this involves division, we'll follow \[d(\frac{ u }{ v }) = \frac{ vdu - udv }{ v ^{2} }\] let us have \[u = x ^{2} ; du = 2x dx\] \[v = e ^{4x} ; dv = e ^{4x} (4) dx\] now you substitute

OpenStudy (anonymous):

i had \[e ^{4x}((2x-4x ^{2})/(e ^{4x}))\]